OpenStudy (anonymous):

What is the vertex and the equation of the axis of symmetry of the graph of y=x^2-6x=7?

OpenStudy (studygurl14):

graph it at www.desmos.com/calulator that should help

OpenStudy (pottersheep):

Do you know how to complete the square?

OpenStudy (pottersheep):

http://www.purplemath.com/modules/sqrvertx.htm :)

OpenStudy (anonymous):

Let's me try to help you. It's about a quadratic function in the form of:y= ax^2 + bx +c The vertex is a point whose x- coordinate value is given by x= -b/2a In your case, b=-6 and a = 1 Once you find the value of x ( you'll find x=3), replace it with x in the function to find the value of y. (in this case, you'll find y=-16). The vertex will be, then (3,-16) The vertical axis passing by x=3 is the axis of symmetry. So the equation of the axis of symmetry is x=3. I hope I have really helped you.

OpenStudy (anonymous):

I have assumed that you meant: y= x^2-6x-7 instead of y=x^2-6x=7

OpenStudy (anonymous):

I'm glad that i could help you. Thank you for the medal!

OpenStudy (studygurl14):

@Jaspharmo I'm the one who gave you the medal, not @txylornichole

OpenStudy (anonymous):

Sorry! I didn't pay attention. Thank you StudyGurl114!

OpenStudy (studygurl14):

lol you're welcome :)